OpenStudy (anonymous):

for f (x) = 4x+1 and g(x) =x^2 - 5 find (f x g) (x) answers 4x^2 - 19 4x^3 + x^2 - 5x - 5 4x^3 + x^2 - 4x - 6 4x^3 + x^2 - 20x - 5

OpenStudy (zzr0ck3r):

(4x+1)(x^2-5)=4x^3-20x+x^2-5=?

OpenStudy (zzr0ck3r):

@dtan5457 \(\times\) not \(\circ\)

OpenStudy (dtan5457):

@zzr0ck3r completely forgot that it isn't all "o"

OpenStudy (anonymous):

So D? @zzr0ck3r

OpenStudy (zzr0ck3r):

correct